A bakery buys sugar from a big distributor to use in baking cakes. Typically, they use 25 pounds of sugar in a day. But depending on the day, they may use a little bit less or more. It is estimated that the standard deviation of demand for sugar is 5 lbs per day. It takes 4 days from the time the bakery orders sugar until the distributor delivers the sugar. The manager wants to have a service level of 0.92. Determine each of the following, assuming that demand is distributed normally.

Example 1.684 for all questions for how it should be answered with decimals

1. Z value for the intended service level?

2. The safety stock that will provide the intended service level?

3. The ROP that will provide the intended service level?
Answer #1

Answer: - Given data

Mean demand = 25 pounds

Standard deviation of demand = 5

Lead time = 4 days

Service level = 0.92 or 92%

a) Z value for given service level = 1.41

b) Safety stock = Z*Std*√L = 1.41*5*√4 = 14.1

c) Reorder point = demad during lead time + safety stock

Demand during lead time = demand per day * lead time

= 25 * 4 = 100 pounds

Reorder point = 100 + 14.1 = 114.1